# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

### v2

@@ II-A Infinite-horizon dynamic games
is the best-response of $p\in\mathcal{P}$ to other players’ plan of action. Under its feedback representation, $\bm{u}^{p}=\bm{K}^{p}\bm{x}\in BR^{p}(\bm{u}^{-p})$ is a solution to the infinite-horizon control problem
@@
$x_{t}\in\mathcal{X},\leavevmode\nobreak\ \leavevmode\nobreak\ u^{p}_{t}\in\mathcal{U}^{p},\leavevmode\nobreak\ \leavevmode\nobreak\ (u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},$ (19c)
@@
While posed in terms of action signals ($\bm{u}^{p}$, $p\in\mathcal{P}$), Problem (19) should be interpreted as the direct search for a best-response policy $\bm{K}^{p}$ against the (fixed) plan of action from other players, $\bm{u}^{-p}\coloneqq(\bm{u}^{\tilde{p}})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$. We slightly abuse notation and let $BR^{p}(\bm{K}^{-p})$ be its solutions, when the problem is parametrised by $\bm{u}^{-p}\coloneqq\bm{K}^{-p}\bm{x}=(\bm{K}^{\tilde{p}}\bm{x})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$. The mapping $BR:\bm{\mathcal{C}}\rightrightarrows\bm{\mathcal{C}}$, defined by $BR(\bm{K})=BR^{1}(\bm{K}^{-1})\times\cdots\times BR^{N_{P}}(\bm{K}^{-N_{P}})$, is the joint best-response to a strategy profile $\bm{K}$. The GFNEs of $\mathcal{G}_{\infty}$ are thus the fixed-points of this mapping: That is, $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}=\{\bm{K}^{*}\in\bm{\mathcal{C}}:\bm{K}^{*}\in BR(\bm{K}^{*})\}$. Due to constraints ($\mathcal{X},\mathcal{U}^{p},\mathcal{U}_{\mathcal{G}}$) and $\bm{\mathcal{C}}^{p}$, an analytical solution to Problem (19) does not exist. Moreover, because infinite-dimensional, its numerical approximation cannot be obtained.
The BRD method for GFNE seeking is given in Algorithm 3. As $\mathcal{G}_{\infty}$ is dynamic and stationary, the procedure needs not episodic repetitions of the game. Instead, the learning dynamics occurs simultaneously with the game’s execution: Players learn and announce their new policies at stages $t\in\{(k{+}1)\Delta T\}_{k\in\mathbb{N}}$. $\bm{K}_{k}\coloneqq(\bm{K}_{k}^{1},\ldots,\bm{K}_{k}^{N_{P}})$ denotes the strategy profile after $k\in\mathbb{N}$ updates. The period $\Delta T\geq 1$ defines the rate at which policies are updated, reflecting some communication structure (e.g., the time needed for each $p\in\mathcal{P}$ to collect $\{\bm{K}_{k}^{\tilde{p}}\}_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$).
@@
Output: GFNE $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$
@@
2 Initialize $\bm{K}_{0}=(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k=0$;
@@ 9            $k\coloneqq k+1$;
If we verbally execute Algorithm 3, we have the following:
@@
The BRD-GFNE induces an operator $T=(1-\eta)I+\eta BR$ which is equivalent to the update rule of its static counterpart. Thus, it possesses the same properties: The learning dynamics converge if $BR$ is non-expansive and the convergence rate is geometric if $BR$ is also a contraction (Theorems 23). These properties can also be stated in terms of stage indices $t\in\mathbb{N}$ by substituting $k=\lfloor t/\Delta T\rfloor$. As in the static case, a careful choice of the learning rate $\eta\in(0,1)$ is required to ensure that this fixed-point iteration is well-defined. Moreover, the BRD-GFNE can be interrupted at any episode $k_{f}>1$ (producing an $\epsilon$-GFNE, $\epsilon>0$) and, if necessary, a central coordinator assigned to project the resulting strategy profile $\bm{K}_{k_{f}}$ onto the feasible set $\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})$. Finally, we note that such considerations can be disregarded when the game is not generalised, that is, when $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$.
III Best-response dynamics via system level synthesis
@@
We focus on $N_{P}$-players linear-quadratic stochastic games $\mathcal{G}_{\infty}^{\text{LQ}}=(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$ with dynamics
@@
defined by matrices $C^{p}\in\mathbb{R}^{N_{z}\times N_{x}}$ and $D^{p\tilde{p}}\in\mathbb{R}^{N_{z}\times N_{u}^{\tilde{p}}}$ with dimension $N_{z}\geq N_{x}+N_{u}$. The following assumptions ensure that GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ are admissible (that is, stabilising):
@@
Finally, the constraint sets $\mathcal{X}$ and $\mathcal{U}^{p}$ ($\forall p$) are convex polyhedra and contain their respective origins: $0\in\mathcal{X}$ and $0\in\mathcal{U}^{p}$.
The class $\mathcal{G}_{\infty}^{\text{LQ}}$ describe problems in which $N_{P}$ non-cooperative agents have to agree on stationary policies that jointly stabilise a global system, robustly to the noise process, while penalising state- and input-deviations differently. While representative of many practically relevant problems, this choice is not restrictive. Our derivations should follow similarly for any collection of cost functions $\{L^{p}\}_{p\in\mathcal{P}}$ satisfying Assumption 2.
III-A System-level best-response mappings
System level synthesis (SLS, [36]) is a novel methodology for controller design which lifts the search for optimal controllers into the search of optimal closed-loop responses to disturbances. Under this approach, the synthesis of stationary control policies can be equivalently posed as the solution to scalable numerical problems, even when operational constraints on the state and control signals need to be enforced. Additionally, the method allows for constraints to be enforced directly on the structure of the control policies. In this section, we present a system-level parametrisation for the best-response mappings in $\mathcal{G}_{\infty}^{\text{LQ}}$.
@@ III-A System-level best-response mappings
$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}^{1}\\ \vdots\\ \bm{\hat{u}}^{N_{P}}\end{bmatrix}$ $=\begin{bmatrix}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\\ \bm{\hat{K}}^{1}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\\ \vdots\\ \bm{\hat{K}}^{N_{P}}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\end{bmatrix}\bm{\hat{w}};$
$=\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}\bm{\hat{w}},$ (23)
where the introduced transfer matrices ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}}$) are denoted as system level responses or closed-loop maps. Under this representation, the following result holds.
Theorem 4 (System level parametrisation, [36]).
@@
    $\begin{bmatrix}zI-A&{-}B^{1}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ {-}B^{N_{P}}\end{bmatrix}\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}=I,$ (24)
@@
    with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$ and $\bm{\tilde{\Phi}_{u}}^{p}=z\bm{\hat{\Phi}_{u}}^{p}$.
@@
To simplify the notation, we momentarily define the matrices $B\coloneqq[B^{1}\ B^{2}\ \cdots\ B^{N_{P}}]$, $\bm{\hat{\Phi}_{u}}\coloneqq\big{[}\bm{\hat{\Phi}_{u}}^{1^{{\mkern-1.5mu\mathsf{T}}}}\leavevmode\nobreak\ \bm{\hat{\Phi}_{u}}^{2^{{\mkern-1.5mu\mathsf{T}}}}\cdots\bm{\hat{\Phi}_{u}}^{N_{P}^{{\mkern-1.5mu\mathsf{T}}}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}$, and $\bm{\hat{K}}\coloneqq\big{[}\bm{\hat{K}}^{1^{\mkern-1.5mu\mathsf{T}}}\leavevmode\nobreak\ \bm{\hat{K}}^{2^{\mkern-1.5mu\mathsf{T}}}\cdots\bm{\hat{K}}^{N_{P}^{\mkern-1.5mu\mathsf{T}}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}$. For the first statement, consider internally stabilising policies $(\bm{\hat{K}}^{1},\ldots,\bm{\hat{K}}^{N_{P}})$ such that $\bm{\hat{u}}^{p}=\bm{\hat{K}}^{p}\bm{\hat{x}}$ ($\forall p\in\mathcal{P}$). Then, from Eq. (23), we have that
$\begin{bmatrix}zI-A&-B\end{bmatrix}\begin{bmatrix}(zI-A-B\bm{\hat{K}})^{-1}\\ \bm{\hat{K}}(zI-A-B\bm{\hat{K}})^{-1}\end{bmatrix}$
$\hskip 15.00002pt=(zI-A)(zI-A-B\bm{\hat{K}})^{-1}{-}B\bm{\hat{K}}(zI-A-B\bm{\hat{K}})^{-1}$
$\hskip 15.00002pt=(zI-A-B\bm{\hat{K}})(zI-A-B\bm{\hat{K}})^{-1}$
$\hskip 15.00002pt=I.$
For the second statement, we first show that $\bm{\hat{K}}$ achieves the desired response, and then that these policies are internally stabilising. Note that Eq. (24) implies that $\bm{\hat{\Phi}_{x}}$ has the leading spectral component $\Phi_{x,1}=I_{N_{x}}$, which is invertible, and thus $\bm{\hat{\Phi}_{x}}^{-1}$ exists. Then, $\bm{\hat{K}}=\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1}$ is well-defined, and we have
$\bm{\hat{x}}$ $=(zI-A-B\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1})^{-1}\bm{\hat{w}}$
$=\bm{\hat{\Phi}_{x}}\left((zI-A)\bm{\hat{\Phi}_{x}}-B\bm{\hat{\Phi}_{u}}\right)^{-1}\bm{\hat{w}}$
$=\bm{\hat{\Phi}_{x}}\bm{\hat{w}},$
due to Eq. (24), and $\bm{\hat{u}}=\bm{\hat{K}}\bm{\hat{x}}=\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1}\bm{\hat{\Phi}_{x}}\bm{\hat{w}}=\bm{\hat{\Phi}_{u}}\bm{\hat{w}}$. Thus, $\bm{\hat{K}}$ achieves the response ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}$), or, equivalently, $(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$ achieve $(\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}})$. To show that this policy is internally stabilising, consider its equivalent representation $\bm{\hat{K}}=\bm{\tilde{\Phi}_{u}}(zI-\bm{\tilde{\Phi}_{x}})^{-1}\bm{\tilde{\Phi}_{y}}$ with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$, $\bm{\tilde{\Phi}_{u}}=z\bm{\hat{\Phi}_{u}}$, and $\bm{\tilde{\Phi}_{y}}=I$, as depicted in Figure 2.
@@
Introducing external perturbations $\{\delta_{x},\delta_{u},\delta_{\xi}\}\subseteq\ell_{\infty}$ into the game, it suffices to verify that the transfer matrices from $(\bm{\hat{\delta}_{x}},\bm{\hat{\delta}_{u}},\bm{\hat{\delta}_{\xi}})$ to $(\bm{\hat{x}},\bm{\hat{u}},\bm{\hat{\xi}})$ are stable. In this case, we have
$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}\\ \bm{\hat{\xi}}\end{bmatrix}=\begin{bmatrix}\bm{\hat{\Phi}_{x}}&\bm{\hat{\Phi}_{x}}B&\bm{\hat{\Phi}_{x}}(zI-A)\\ \bm{\hat{\Phi}_{u}}&I+\bm{\hat{\Phi}_{u}}B&\bm{\hat{\Phi}_{u}}(zI-A)\\ \frac{1}{z}I&\frac{1}{z}B&\frac{1}{z}(zI-A)\\ \end{bmatrix}\begin{bmatrix}\bm{\hat{\delta}_{x}}\\ \bm{\hat{\delta}_{u}}\\ \bm{\hat{\delta}_{\xi}}\end{bmatrix}.$ (26)
Since $\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}\in\frac{1}{z}\mathcal{RH}_{\infty}$, the transfer matrices in Eq. (26) are all stable. We thus conclude that the policy $\bm{\hat{K}}$ is internally stabilising, and so are the policies $(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$. ∎
@@
A policy $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($\forall p\in\mathcal{P}$) is defined by the kernel $\bm{\Phi}^{p}=\bm{\Phi_{u}}^{p}*\bm{\Phi_{x}}^{-1}$, and implemented as
@@
The system level parametrisation enables a methodology for policy synthesis consisting of searching the space of stabilising policies (in)directly through $(\bm{\Phi_{x}},\bm{\Phi_{u}}^{p})$, $p\in\mathcal{P}$. In particular, this parametrisation can be leveraged to reformulate the best-response dynamics maps in $\mathcal{G}_{\infty}^{\text{LQ}}$ as tractable numerical programs. In this direction, consider that players design stabilising policies $\bm{K}=(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})$ by choosing their desired system level responses $\bm{\Phi_{u}}=(\bm{\Phi_{u}}^{1},\ldots,\bm{\Phi_{u}}^{N_{P}})$ simultaneously. From the affine space Eq. (24), the signal $\bm{\Phi_{x}}$, common to all players, satisfies the (deterministic) linear dynamics
@@
$\bm{F_{\Phi}}:\bm{\Phi_{u}}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\displaystyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{\Phi_{u}}^{p}+\bm{\delta}I_{N_{x}}\Big{)},$ (29)
Using Parseval’s Theorem [42], we substitute Eq. (23) into Eq. (21) to redefine the objective functionals of $\mathcal{G}_{\infty}^{\text{LQ}}$ as
$J^{p}(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})$
$\quad=\mathrm{E}\Big{[}\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}w_{n}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}w_{n}\|_{F}^{2}\Big{)}\Big{]}$
$\quad=\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\|_{F}^{2}\Big{)}.$
@@
$U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})\coloneqq\{\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}:(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})\in\bm{F_{\Phi}}^{\dagger}(\bm{\mathcal{C}}_{x}),\\ \bm{\Phi_{u}}^{p}*\bm{w}\in U^{p}(\bm{\Phi_{u}}^{-p}*\bm{w})\},$

### v3

@@ 2.1 Infinite-horizon dynamic games
is the best-response of $p\in\mathcal{P}$ to other players’ plan of action. Under its feedback representation, $\bm{u}^{p}=\bm{K}^{p}\bm{x}\in BR^{p}(\bm{u}^{-p})$ is a solution to the infinite-horizon stochastic control problem
@@
$x_{t}\in\mathcal{X},~{}~{}u^{p}_{t}\in\mathcal{U}^{p},~{}~{}(u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},$ (17c)
@@
While posed in terms of action signals ($\bm{u}^{p}$, $p\in\mathcal{P}$), Problem (17) should be interpreted as the direct search for a best-response policy $\bm{K}^{p}$ against the (fixed) plan of action from other players, $\bm{u}^{-p}\coloneqq(\bm{u}^{\tilde{p}})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$. We slightly abuse notation and let $BR^{p}(\bm{K}^{-p})$ be its solutions when parametrised by $\bm{u}^{-p}\coloneqq\bm{K}^{-p}\bm{x}=(\bm{K}^{\tilde{p}}\bm{x})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$. The mapping $BR:\bm{\mathcal{C}}\rightrightarrows\bm{\mathcal{C}}$, defined by $BR(\bm{K})=BR^{1}(\bm{K}^{-1})\times\cdots\times BR^{N_{P}}(\bm{K}^{-N_{P}})$, is the joint best-response to a strategy profile $\bm{K}\in\bm{\mathcal{C}}$. The GFNE of $\mathcal{G}_{\infty}$ thus correspond to the fixed-points of this mapping: That is, $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}=\{\bm{K}^{\star}\in\bm{\mathcal{C}}:\bm{K}^{\star}\in BR(\bm{K}^{\star})\}$. Due to constraints ($\mathcal{X},\mathcal{U}^{p},\mathcal{U}_{\mathcal{G}}$) and $\bm{\mathcal{C}}^{p}$, an analytical solution to Problem (17) does not exist. Moreover, because infinite-dimensional, its numerical approximation cannot be obtained.
A BRD for GFNE seeking is outlined in Algorithm 3. As $\mathcal{G}_{\infty}$ is dynamic and stationary, the procedure does not require episodic repetitions of the game. Instead, the learning dynamics occurs simultaneously with the game’s execution: Players learn and announce their new policies at stages $t\in\{(k{+}1)\Delta T\}_{k\in\mathbb{N}}$. $\bm{K}_{k}\coloneqq(\bm{K}_{k}^{1},\ldots,\bm{K}_{k}^{N_{P}})$ denotes the strategy profile after $k\in\mathbb{N}$ updates. The period $\Delta T\geq 1$ defines the rate at which policies are updated, reflecting some communication structure (e.g., the time needed for each $p\in\mathcal{P}$ to collect $\{\bm{K}_{k}^{\tilde{p}}\}_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$).
@@
Output: GFNE $\bm{K}^{\star}=(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$
@@
2 Initialize $\bm{K}_{0}\coloneqq(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k\coloneqq 0$;
@@ 9            $k\coloneqq k+1$;
A verbal execution of Algorithm 3 yields the following:
@@
The BRD-GFNE induces an operator $T=(1-\eta)I+\eta BR$ which is equivalent to the update rule of its static counterpart. Thus, it possesses the same properties: The learning dynamics converge if $BR$ is nonexpansive and the convergence rate is geometric if $BR$ is also a contraction (Lemmas 12). These properties can also be stated in terms of stage indices $t\in\mathbb{N}$ by replacing $k=\lfloor t/\Delta T\rfloor$. As in the static case, a careful choice of the learning rate $\eta\in(0,1)$ is required to ensure that this fixed-point iteration is well-defined. Finally, we stress that the BRD-GFNE can be interrupted at any $k_{f}>1$, thus producing an $\epsilon$-GFNE policy $\bm{K}_{k_{f}}$ with associated equilibrium gap $\epsilon>0$.
3 Best-response dynamics via system level synthesis
@@
We focus on $N_{P}$-player linear-quadratic stochastic games $\mathcal{G}_{\infty}^{\text{LQ}}=(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$ with dynamics
@@
defined by matrices $C^{p}\in\mathbb{R}^{N_{z}\times N_{x}}$ and $D^{p\tilde{p}}\in\mathbb{R}^{N_{z}\times N_{u}^{\tilde{p}}}$ with dimension $N_{z}\geq N_{x}+N_{u}$. The following assumptions ensure that stabilising GFNE solutions to $\mathcal{G}_{\infty}^{\text{LQ}}$ exist:
@@
Finally, the sets $\mathcal{X}$, $\mathcal{U}^{p}$ ($\forall p$), and $\mathcal{U}_{\mathcal{G}}$, are convex polyhedra satisfying $0\in\mathbf{relint}~{}\mathcal{X}$, $0\in\mathbf{relint}~{}\mathcal{U}^{p}$, and $0\in\mathbf{relint}~{}\mathcal{U}_{\mathcal{G}}$.
The class $\mathcal{G}_{\infty}^{\text{LQ}}$ describe problems in which $N_{P}$ noncooperative agents have to agree on stationary policies that jointly stabilise a global system, robustly to the noise process, while penalising state- and input-deviations differently. While representative of many practically relevant problems, this choice is not restrictive. Our derivations should follow similarly for any collection of cost functions $\{L^{p}\}_{p\in\mathcal{P}}$ satisfying Assumption 2.
3.1 System-level best-response mappings
System level synthesis (SLS, [35]) is a novel methodology for controller design centred on the equivalent representation of control policies in terms of the closed-loop responses that they achieve. Unlike similar approaches, such as the Youla [40] and input-output (IOP, [41]) parametrisations, SLS allows the synthesis of state-feedback policies to be posed as the solution to convex optimisation problems, even when subjected to constraints on the state- and input-signals, and on the structure of the policy itself. In this section, we present a system-level parametrisation for the best-response mappings in $\mathcal{G}_{\infty}^{\text{LQ}}$.
@@ 3.1 System-level best-response mappings
$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}^{1}\\ \vdots\\ \bm{\hat{u}}^{N_{P}}\end{bmatrix}$ $=\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}\bm{\hat{w}},$ (21)
where $\bm{\hat{\Phi}}_{x}=(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}$ and $\bm{\hat{\Phi}}_{u}^{p}=\bm{\hat{K}}^{p}\bm{\hat{\Phi}}_{x}$ ($p\in\mathcal{P}$). The introduced transfer matrices ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}}$) are referred to as system level responses or closed-loop maps. Under this representation, the following result holds.
Theorem 1 (System level parametrisation).
@@
    $\begin{bmatrix}zI-A&{-}B^{1}~{}\cdots~{}{-}B^{N_{P}}\end{bmatrix}\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}=I,$ (22)
@@
    with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$ and $\bm{\tilde{\Phi}_{u}}^{p}=z\bm{\hat{\Phi}_{u}}^{p}$ (see Figure 2).
@@
Defining $B\coloneqq[B^{1}\ B^{2}\ \cdots\ B^{N_{P}}]$, and transfer matrices $\bm{\hat{\Phi}_{u}}\coloneqq\texttt{col}(\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}})$ and $\bm{\hat{K}}\coloneqq\texttt{col}(\bm{\hat{K}}^{1},\ldots,\bm{\hat{K}}^{N_{P}})$, the proof is as in [35, Theorem 4.1]. We refer to the Supplementary Material for the full details. In the second statement, we consider an alternative representation $\bm{\hat{K}}=\bm{\tilde{\Phi}_{u}}(zI-\bm{\tilde{\Phi}_{x}})^{-1}\bm{\tilde{\Phi}_{y}}$ with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$, $\bm{\tilde{\Phi}_{u}}=z\bm{\hat{\Phi}_{u}}$, and $\bm{\tilde{\Phi}_{y}}=I$: This leads to the transfer matrices from $(\bm{\hat{\delta}_{x}},\bm{\hat{\delta}_{u}},\bm{\hat{\delta}_{\xi}})$ to $(\bm{\hat{x}},\bm{\hat{u}},\bm{\hat{\xi}})$,
$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}\\ \bm{\hat{\xi}}\end{bmatrix}=\begin{bmatrix}\bm{\hat{\Phi}_{x}}&\bm{\hat{\Phi}_{x}}B&\bm{\hat{\Phi}_{x}}(zI-A)\\ \bm{\hat{\Phi}_{u}}&I+\bm{\hat{\Phi}_{u}}B&\bm{\hat{\Phi}_{u}}(zI-A)\\ \frac{1}{z}I&\frac{1}{z}B&\frac{1}{z}(zI-A)\\ \end{bmatrix}\begin{bmatrix}\bm{\hat{\delta}_{x}}\\ \bm{\hat{\delta}_{u}}\\ \bm{\hat{\delta}_{\xi}}\end{bmatrix},$ (24)
which are all stable due to $\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}\in\frac{1}{z}\mathcal{RH}_{\infty}$. Thus, the policy $\bm{\hat{K}}=(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$ is internally stabilising.∎
@@
A policy $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($p\in\mathcal{P}$) is defined by the kernel $\bm{\Phi}^{p}=\bm{\Phi_{u}}^{p}*\bm{\Phi_{x}}^{-1}$, and can be implemented as
@@
The system level parametrisation enables a methodology for policy synthesis consisting of searching the space of stabilising policies (in)directly through $(\bm{\Phi_{x}},\bm{\Phi_{u}}^{p})$, $p\in\mathcal{P}$. In particular, this parametrisation can be leveraged to reformulate the best-response mappings in $\mathcal{G}_{\infty}^{\text{LQ}}$ as numerically tractable problems. In this direction, consider that players design stabilising policies $\bm{K}=(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})$ by choosing their desired system level responses $\bm{\Phi_{u}}=(\bm{\Phi_{u}}^{1},\ldots,\bm{\Phi_{u}}^{N_{P}})$, simultaneously. From the affine space Eq. (22), the signal $\bm{\Phi_{x}}$, common to all players, satisfies the (deterministic) linear dynamics
@@
$\bm{F_{\Phi}}:\bm{\Phi_{u}}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\textstyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{\Phi_{u}}^{p}+\bm{\delta}I_{N_{x}}\Big{)}.$ (27)
Using the system level responses, the objective functionals of $\mathcal{G}_{\infty}^{\text{LQ}}$ can be shown as equivalent to the functional
$J^{p}(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})\\ =\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\|_{F}^{2}\Big{)}.$
@@
$U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})\coloneqq\{\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}:\bm{F_{\Phi}}\bm{\Phi_{u}}\in\bm{\mathcal{C}}_{x},\\ \quad\bm{\Phi_{u}}^{p}*\bm{w}\in U^{p}(\bm{\Phi_{u}}^{-p}*\bm{w})\},$